OpenStudy (kelgar2006):

Paul took 900 US dollars to the bank to exchange for Euros. The exchange rate on that day was 0.75 Euros per U.S. dollar. How many Euros did he get in exchange for 900 US dollars.

OpenStudy (mathstudent55):

Look at the exchange rate: 0.75 Euros = $1 This can be written as a fraction two ways: \(\dfrac{0.75~Euro}{$1} = \dfrac{$1}{0.75~Euro} = 1\) Do you follow so far?

OpenStudy (mathstudent55):

Now start with the amount you need to convert: \($900\) You have US dollars, but you want Euros. That means you need to multiply $900 by a conversion fraction that has US Dollars in the denominator (so the dollars will cancel out) and Euros in the numerator, so your answer will be in Euros. \($900 \times \dfrac{0.75~Euros}{$1} \) The dollars cancel out, and you are left with Euros.

OpenStudy (mathmale):

kelgar: would you please "do the math" here? You final answer should be in Euros.